OpenStudy (anonymous):

What is the solution to the equation? t(9 – 1) = 5 • 9 – 5 • 1 A. –1 B. 1 C. 5 D. 9

OpenStudy (anonymous):

Distribute the t on the left 9t-t=5*9-5*1 multiplication first, then subtraction 8t=45-5 8t=40 t=5